# 1D Transition Metal Oxide Chains as a Challenging Model for Ab Initio Calculations

### v1

@@ III.3 Electronic properties
The study of transition metal oxides (TMOs) is challenging due to the presence of localized 3d orbitals in transition metal atoms. The PBE approximation suffers from self-interaction errors, which prevent it from accurately describing these localized orbitals. This limitation is evident in the PBE band structures of the FM state of 1D-TMOs, where only MnO exhibits a band gap (see Fig. 5). To correctly open the band gap, PBE must be supplemented with correction methods such as DFT+$U$.
The converged $U$ values, calculated at the optimal lattice parameter, are presented in Table 2.
@@
For VO, the total energy varies significantly with different lattice parameters when using the computed $U$ value from DFPT calculations. Unfortunately, we were unable to resolve this issue, so we adopted a fixed $U$ value of 4.8 eV for V Nakamura et al. (2006).
Figure 5: Calculated band structure of all considered systems in FM magnetic state. The Fermi level $E_{F}$ is chosen as the zero of the energy scale. Effective values of $U$ as determined using QE package in Table.2 are used in the DFT+$U$ calculation. Top panels shows the PBE band structures and the bottom panels shows the DFT+$U$ band structures. For VO and CrO, the $U$ parameters are set to 4.8 eV and 5.0 eV, respectively Nakamura et al. (2006). The red and blue lines represent the spin-up (up) and spin-down (dn) band structures, respectively.
Figure 6: Calculated band structure of all considered systems in AFM magnetic state. The Fermi level $E_{F}$ is chosen as the zero of the energy scale. Effective values of U as determined using QE package in Table.2 are used in the DFT+$U$ calculation. Top panels shows the PBE band structures and the bottom panels shows the DFT+$U$ band structures. For VO and CrO, the $U$ parameters are set to 4.8 eV and 5.0 eV, respectively Nakamura et al. (2006).
Table 2: Converged value of the effective Hubbard parameter $U$ (in eV) for the transition metal(TM) 3d for FM and AFM states.
@@
Figures 5 and 6 show the band structures of the 1D-TMOs computed using the PBE (top panel) and DFT+$U$(bottom panel) methods for AFM and FM magnetic states, respectively.
For the FM states of VO, FeO, and CoO, the PBE method predicts a half-metallic character, while it predicts a fully metallic state for NiO and a semi-metallic state for CrO. Applying DFT+$U$ results in an insulating state for all 1D-TMOs. MnO is the only case where both PBE and DFT+$U$ consistently predict an insulating state, making it a straightforward candidate for further ab initio studies.
@@
In the AFM state (Fig.  6), both the PBE and DFT+$U$ methods predict an insulating ground state for VO, MnO, FeO, and NiO. The $U$ correction pushes the occupied bands down and opens the energy gap. In the case of CrO and CoO, PBE fails to open a gap, predicting a metallic ground state. However, the $U$ correction succeeds in predicting insulating behavior.
@@
For CrO in the FM state, a straightforward application of DFT+$U$ does not produce an insulating solution. To assess whether the FM state is inherently semi-metallic or if this result arises from the system being trapped in a local minimum, it is important to test various initializations of the $d$-orbital density matrix within DFT+$U$Dorado et al. (2009, 2013). A simpler and more effective alternative is to randomly initialize the magnetic moments on the oxygen atomsPayami Shabestar (2022). Using this approach, we find that the FM state of CrO can indeed transition to an insulating state.
III.4 Comparison between DFT+$U$ and CCSD
@@
CCSD is known for its high accuracy in electronic structure predictions, but it comes with a significantly higher computational expense compared to other methods, such as DFT+$U$, due to its steep scaling with system size. Due to having less number of electrons in 1D-TMOs, they can be ideal test cases to compare DFT+$U$ result with CCSD.
@@
The total energy versus lattice parameter for 1D-TMOs is calculated using CCSD (Fig. 4). The calculated energy differences, $\Delta E$, (in meV) for CrO, MnO, FeO, CoO and NiO chains using DFT+$U$ and CCSD methods (starting from UHF orbitals) are presented in Table 1. Notably, the CCSD method generally predicts higher energy differences compared to DFT+$U$. For CrO, MnO, CoO and NiO, the CCSD results show significantly higher values for $\Delta E$ with respect to the DFT+$U$ ones. In contrast, FeO demonstrates closer agreement between the two approaches, with only a slight deviation in the energy difference.
@@
The largest discrepancies are observed for NiO and CrO. For NiO, DFT+$U$ predicts an energy difference of $\Delta E=-230$ meV, while CCSD yields a much lower value of $\Delta E=-596$ meV, highlighting the challenges of accurately capturing electron correlation in this material. For CrO, the discrepancy is even more striking: DFT+$U$ predicts the ferromagnetic (FM) state as the ground state with $\Delta E=+128$ meV, whereas CCSD identifies the antiferromagnetic (AFM) state as the true ground state with $\Delta E=-215$ meV.
@@
The discrepancy between PBE/DFT+$U$ and CCSD in predicting the ground state of CrO may partly stem from the use of unrestricted Hartree-Fock (UHF) as the reference for CCSD calculations, as UHF itself favors the antiferromagnetic (AFM) state as the ground state of CrO. Since CCSD builds upon UHF wavefunctions, this preference may influence the final result. We find that both FHI-aims and PySCF predict the AFM ground state for CrO at the UHF level, suggesting that this outcome is likely robust with respect to the choice of basis set.
@@
The equilibrium lattice constants of the CCSD closely match the PBE results, whereas DFT+$U$ predicts lattice constants that are 3–4% larger. This discrepancy highlights the impact of the Hubbard $U$ correction when combined with a GGA functional like PBE, which tends to overestimate lattice constants–a well-known issue in DFT+$U$ calculations (see Fig. 4 and Table 1).
@@ IV Conclusion
In summary, we investigated the structural, electronic, and magnetic properties of one-dimensional transition-metal oxide chains using a range of ab-initio methods. We began by addressing convergence and instability issues that arose in our DFT calculations, particularly for CoO and NiO, where wave function instability resulted in convergence to wrong states rather than ground states. Using stable wave functions at some lattice parameters as the starting point, we were able to overcome these challenges and achieve stable ground-state configurations. For ground state properties, we compared results from pseudopotential (Quantum ESPRESSO) and full-potential (FHI-aims) methods. The results showed general agreement in lattice constants and energy differences.
@@
Lastly, a comparison between DFT+$U$ and the more accurate, yet computationally demanding, CCSD method revealed significant discrepancies, particularly for NiO and CrO. In the case of CrO, we encountered a notable contradiction: CCSD predicts an AFM ground state, whereas DFT+$U$ favors an FM ground state. This makes CrO chain a compelling candidate for benchmarking and testing the accuracy of various ab initio methods.
For all cases, the CCSD method predicts significantly larger energy differences between AFM and FM states compared to DFT+$U$, suggesting that the $U$ parameters obtained through linear response theory may be overestimated for these systems. However, for a more reliable comparison, CCSD calculations should be performed using larger basis sets. Overall, this comprehensive benchmarking highlights the strengths and limitations of each method, providing valuable insights for future research on low-dimensional transition-metal oxides, which serve as a challenging model for ab initio calculations.
@@ Acknowledgments
SdG work was supported by the European Commission through the MAX Centre of Excellence for supercomputing applications (grant numbers 10109337 and 824143) and by the Italian MUR, through the Italian National Centre from HPC, Big Data, and Quantum Computing (ICSC, grant number CN00000013). Computational resources were provided by CINECA.

### v2

@@ III.3 Electronic properties
The study of transition metal oxides (TMOs) is challenging due to the presence of localized 3dd orbitals in transition metal atoms. The PBE approximation suffers from self-interaction errors, which prevent it from accurately describing these localized orbitals. This limitation is evident in the PBE band structures of the FM state of 1D-TMOs, where only MnO exhibits a band gap (see Fig. 5). To correctly open the band gap, PBE must be supplemented with correction methods such as DFT+$U$. The converged $U$ values, calculated at the optimal lattice parameter, are presented in Table. 2.
@@
For VO, we encountered serious convergence problems when trying to compute the Hubbard $U$ value self-consistently using DFPT. The SCF calculations were unstable across different lattice parameters and resulted in large, unphysical fluctuations in total energy. Because of this, we were unable to obtain a reliable $U$ value and instead used a fixed value of 4.8 eV for V Nakamura et al. (2006). This approach ensured numerical stability and produced results that are consistent with the general behavior observed in other TMOs.
Figure 5: Calculated band structure of all considered systems in the FM magnetic state. The Fermi level $E_{F}$ is chosen as the zero of the energy scale. Effective values of $U$ as determined using the QE package in Table. 2 are used in the DFT+$U$ calculation. Top panels show the PBE band structures and the bottom panels show the DFT+$U$ band structures. For VO, the $U$ parameteris set to 4.8 eV Nakamura et al. (2006). The red and blue lines represent the spin-up (up) and spin-down (dn) band structures, respectively.
Figure 6: Calculated band structure of all considered systems in the AFM magnetic state. The Fermi level $E_{F}$ is chosen as the zero of the energy scale. Effective values of $U$ as determined using the QE package in Table. 2 are used in the DFT+$U$ calculation. Top panels show the PBE band structures and the bottom panels show the DFT+$U$ band structures. For VO, the $U$ parameter is set to 4.8 eV Nakamura et al. (2006).
Table 2: Converged value of the effective Hubbard $U$ parameter (in eV) for the transition metal(TM) 3d orbitals for FM and AFM states.
@@
Fig. 5 and  6 show the band structures of the 1D-TMOs computed using the PBE (top panel) and DFT+UU(bottom panel) methods for AFM and FM magnetic states, respectively.
For the FM states of VO, FeO, and CoO, the PBE method predicts a half-metallic character, while it predicts a fully metallic state for NiO and a semi-metallic state for CrO. Applying DFT+$U$ results in an insulating state for all 1D-TMOs. For NiO, while our PBE calculations predict a metallic ferromagnetic state, applying DFT+$U$ opens a band gap. In a study on an ideal infinite NiO chain using a hybrid functional Jacob and Palacios (2006), a metastable half-metallic solution was obtained for the ferromagnetic ordering, underscoring the importance of carefully identifying the global minimum in the electronic structure of 1D-TMOs.
Among the TMOs we studied, MnO is the only system that exhibited stable convergence across all PBE and DFT+$U$ calculations, regardless of the code used, and showed no signs of spurious local minima or metastable solutions. In addition, both PBE and DFT+$U$ consistently predict an insulating ground state for MnO in our 1D chain calculations. This is consistent with bulk studies, which also report an insulating character at the PBE level—typically yielding a gap of  1.1eV—and larger values ( 3.5eV) from more accurate methods such as CCSD[ Gao et al. (2020); Pask et al. (2001)]. This combination of numerical stability and robust insulating behavior makes MnO a particularly suitable candidate for benchmarking correlated electronic structure methods across different dimensionalities.
@@
In the AFM state (Fig. 6), both the PBE and DFT+$U$ methods predict an insulating ground state for VO, MnO, FeO, and NiO. The $U$ correction pushes the occupied bands down and opens the energy gap. In the case of CrO and CoO, PBE fails to open a gap, predicting a metallic ground state. However, the $U$ correction succeeds in predicting insulating behavior.
@@
For CrO in the FM state, a straightforward application of DFT+$U$ does not produce an insulating solution. To assess whether the FM state is inherently semi-metallic or if this result arises from the system being trapped in a local minimum, it is important to test various initializations of the $d$-orbital density matrix within DFT+UUDorado et al. (2009, 2013). A simpler and more effective alternative is to randomly initialize the magnetic moments on the oxygen atoms Payami (2021). Using this approach, we find that the FM state of CrO can indeed transition to an insulating state.
Figure 7: Trend of (a) local magnetic moments (M) in $\mu_{B}$ and (b) band gaps (Eg) in eV for selected transition metal monoxides (MnO, FeO, CoO, NiO) as obtained for 1D-TMOs and from reported DFT+$U$ studies on bulk systems. In panel (a), the data points are slightly shifted along the $x$-axis for clarity.
Fig. 7(a) compares our calculated magnetic moments for MnO, FeO, CoO, and NiO with those reported in previous DFT+$U$ studies of bulk systems[ Anisimov et al. (1991); Qu et al. (2022); Trimarchi et al. (2018); Schrön et al. (2012)]. Due to low coordination, magnetic moments in 1D-TMOs are generally expected to be larger than in bulk. While our results qualitatively follow bulk trends, the magnitudes are not consistently higher. These deviations are primarily due to differences in computational methodology. In particular, definitions of local moments vary across studies—e.g., via atomic sphere integration or orbital projections. Here, we extract magnetic moments from the PDOS.
Fig. 7(b) shows that the calculated band gaps in the 1D-TMO chains are systematically larger than those in bulk[ Anisimov et al. (1991); Qu et al. (2022); Trimarchi et al. (2018)], consistent with enhanced electron localization in reduced dimensions. For all four compounds, the 1D gaps exceed bulk DFT+$U$ values by 0.3–1.2eV. While this reflects stronger correlation effects in 1D, the precise values remain sensitive to the choice of $U$, and the exchange-correlation functional.
Overall, the observed trends in both magnetic moments and band gaps indicate stronger localization and enhanced correlation effects in the 1D models compared to their bulk counterparts. This systematic behavior underscores the utility of 1D-TMOs as simplified yet physically meaningful platforms for benchmarking electronic structure methods in strongly correlated systems.
III.4 Comparison between DFT+UUnd CCSD
@@
CCSD is known for its high accuracy in electronic structure predictions, but it comes with a significantly higher computational expense compared to other methods, such as DFT+$U$, due to its steep scaling with system size. Due to having fewer of electrons in 1D-TMOs, they can be ideal test cases to compare DFT+$U$ results with CCSD.
@@
The total energy vs lattice parameter for 1D-TMOs is calculated using CCSD (Fig. 4). The calculated energy differences, $\Delta E$, (in meV) for CrO, MnO, FeO, CoO, and NiO chains using DFT+$U$ and CCSD methods (starting from UHF orbitals) are presented in Table. 1. Notably, the CCSD method generally predicts higher energy differences compared to DFT+$U$. For CrO, MnO, CoO, and NiO, the CCSD results show significantly higher values for $\Delta E$ with respect to the DFT+$U$. In contrast, FeO demonstrates closer agreement between the two approaches, with only a slight deviation in the energy difference.
@@
The largest discrepancies are observed for NiO and CrO. For NiO, DFT+$U$ predicts an energy difference of $\Delta E=-230$ meV, while CCSD yields a much lower value of $\Delta E=-596$ meV, highlighting the challenges of accurately capturing electron correlation in this material. For CrO, the discrepancy is even more striking: DFT+$U$ predicts the ferromagnetic state as the ground state with $\Delta E=+128$ meV, whereas CCSD identifies the antiferromagnetic state as the true ground state with $\Delta E=-215$ meV.
@@
The discrepancy between PBE, DFT+$U$ and CCSD in predicting the ground state of CrO may partly stem from the use of unrestricted Hartree-Fock (UHF) as the reference for CCSD calculations, as UHF itself favors the antiferromagnetic state as the ground state of CrO. Since CCSD builds on UHF reference wavefunctions, this preference may influence the final result. We find that both FHI-aims and PySCF predict the AFM ground state for CrO at the UHF level, suggesting that this outcome is likely robust with respect to the choice of basis set.
@@
The equilibrium lattice constants of the CCSD closely match the PBE results, whereas DFT+$U$ predicts lattice constants that are 3%–4% larger. This discrepancy highlights the impact of the Hubbard $U$ correction when combined with a GGA functional like PBE, which tends to overestimate lattice constants—a well-known issue in DFT+$U$ calculations (see Fig. 4 and Table. 1).
@@ IV CONCLUSION
In summary, we investigated the structural, electronic, and magnetic properties of one-dimensional transition-metal oxide chains using a range of ab initio methods. We began by addressing convergence and instability issues that arose in our DFT calculations, particularly for CoO, and NiO, where wavefunction instability resulted in convergence to incorrect states rather than ground states. Using stable wavefunctions at some lattice parameters as the starting point, we were able to overcome these challenges and achieve stable ground-state configurations. For ground state properties, we compared results from pseudopotential (QE) and full-potential (FHI-aims) methods. The results showed general agreement in lattice constants and energy differences.
@@
Finally, a comparison between DFT+$U$ and the more accurate, yet computationally demanding, CCSD method revealed significant discrepancies, particularly for NiO and CrO. In the case of CrO, we encountered a notable contradiction: CCSD predicts an AFM ground state, whereas DFT+$U$ favors an FM ground state. This makes the CrO chain a compelling candidate for benchmarking and testing the accuracy of various ab initio methods.
For all cases, the CCSD method predicts significantly larger energy differences between AFM and FM states compared to DFT+$U$, suggesting that the $U$ parameters obtained through the linear response theory may be overestimated for these systems. However, for a more reliable comparison, CCSD calculations should be performed using larger basis sets. Comparison with bulk TMO data reveals that 1D models capture key trends, including the decrease in magnetic moment and the increase in band gap with $d$-orbital filling. Magnetic moments generally follow bulk behavior, while band gaps are consistently larger in 1D systems due to enhanced electron localization from reduced dimensionality. These findings support 1D-TMOs as simplified yet informative benchmarks for correlated electronic structure methods. Overall, this comprehensive benchmarking highlights the strengths and limitations of each method, providing valuable insights for future research on low-dimensional transition-metal oxides, which serve as a challenging model for ab initio calculations.
DATA AVAILABILITY
The data that support the findings of this study are available upon reasonable request from the authors.
@@ ACKNOWLEDGMENTS
This work was supported by the Vice-Chancellor for Research Affairs of Isfahan University of Technology (IUT). S.d.G.’s work was supported by the European Commission through the MAX Center of Excellence for Supercomputing Applications (Grant Nos. 10109337 and 824143) and by the Italian MUR, through the Italian National Center for HPC, Big Data, and Quantum Computing (ICSC, Grant No. CN00000013). Computational resources were provided by CINECA.
APPENDIX: HUBBARD $U$ PARAMETERS COMPUTED FROM DFPT
In this appendix, we provide the computed Hubbard $U$ parameters obtained via density-functional perturbation theory (DFPT) for each transition metal oxide (TMO) chain considered in this work. These data correspond to the values used in all DFT+$U$ total energy and electronic structure calculations presented in the main text.
As explained in Sec. II in our DFT+$U$ calculations, the Hubbard $U$ parameter was not treated as an empirical input but instead was computed self-consistently using the linear response approach within DFPT, following the method proposed by Cococcioni and de Gironcoli. Cococcioni and De Gironcoli (2005), This technique allows for a first-principles determination of $U$ by evaluating the response of the localized $d$-orbitals to a perturbing potential within the same theoretical framework as the DFT functional.
For each material (CrO, MnO, FeO, CoO, and NiO), the $U$ parameter was computed as a function of the lattice constant separately for the FM and AFM phases. These values were used consistently in our PBE+$U$ calculations for total energies and electronic structures. Fig. 8 summarizes the computed Hubbard $U$ values across the lattice scans for all systems.
As shown, the $U$ parameter generally decreases with increasing lattice constant, likely due to changes in TM–O hybridization that enhance electronic screening in the expanded structures. In addition, the observed differences between FM and AFM phases indicate a non-negligible dependence of $U$ on the magnetic configuration. These variations are important when computing relative phase stability, as they influence the total energy landscape in DFT+$U$.
Figure 8: Hubbard $U$ values calculated via DFPT as a function of lattice constant(d) for each TMO chain in both FM and AFM magnetic configurations.